if in a triangle
∠a = 2∠b = 6∠c
find ∠a, ∠b, ∠c
please help

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Hint: A + B + C = 180 degrees

Answer by Edwin McCravy(20055)


Hey, don't use < for an angle for if you put anything just to the
right of it it will delete it.  Use 6 & # 8 7 3 6 ; (without the spaces) instead.


∠a = 2∠b = 6∠c

Get ∠a and ∠b in terms of ∠c

∠a = 6∠c   <--- that's ∠a in terms of ∠c

2∠b = 6∠c

so

∠b = 3∠c   <--- that's ∠b in terms of ∠c


Substitute in

  ∠a + ∠b + ∠c = 180°

6∠c + 3∠c + ∠c = 180°

           10∠c = 180°

Divide both sides by 10

             ∠c = 18°

and since ∠b = 3∠c

            ∠b = 3(18°)
            ∠b = 54°

and since ∠a = 6∠c

            ∠a = 6(18°)
            ∠b = 108°

Edwin